OpenStudy (anonymous):

Dont know how to do this! If AB = 6.5, CD = 8, and FE = 22.5, calculate the perimeter of Triangle AEC

OpenStudy (anonymous):

its hard no figure

OpenStudy (anonymous):

P = a + b + c P = 6.5 + 8 + 22.5 P = 37

OpenStudy (phi):

you should draw or post the figure. Where are points A,B,C,D,E,F??